OpenStudy (anonymous):

1/5 times radical x-3

OpenStudy (nfcfox):

Does x have a value?

OpenStudy (anonymous):

nope

OpenStudy (nfcfox):

What is the question asking?

OpenStudy (anonymous):

its just the 1/5 is right next to the radical and under the radical it says x-3. so its like y=(1/5)radical and under that is x-3

OpenStudy (nfcfox):

Ok? What is it asking though.

OpenStudy (anonymous):

oh lol sorry its asking l ( x ) = 15 x − 3 i just wanted to solve it for the inverse

OpenStudy (anonymous):

and to see if it is a function or not

OpenStudy (nfcfox):

So it gave 15x-3 and is asking for the inverse?

OpenStudy (anonymous):

this was the original problem ( x ) = 15 x − 3 so i just did the x and y switch thing to set it up as x= and then i was gonna solve for y. but when i asked it on here i just set it up as any equation i apologize.

OpenStudy (anonymous):

15x-3 it doesn't like when i type that sorry lol

OpenStudy (anonymous):

wait it says 1/5 and the x-3 is under a radical

OpenStudy (anonymous):

this problem is giving me problems lol!

OpenStudy (anonymous):

so its (x)= (1/5) times radical x-3